OpenStudy (anonymous):

A skateboarder has an accerlation of -1.9m/s^2 if her initial speed is 6 m/s how long does it take her to stop?

OpenStudy (anonymous):

AS, v=u+at ,', 0 = 6 -1.9t {v=0, as, finally it'll stop} =>1.9t=6 =>t=6\1.9=3.16 sec...